OpenStudy (oswaldmurphy):

Write the equation of the line that passes through the points (3, 6) and (4, 10) using function notation. f(x) = 4x − 6 f(x) = x + 4 y = x + 4 y = 4x − 6

OpenStudy (oswaldmurphy):

@pooja195

OpenStudy (oswaldmurphy):

@mathstudent55

OpenStudy (archaists):

@OswaldMurphy it's f(x) = 4x − 6

OpenStudy (oswaldmurphy):

Is it because its crossing and it goes through x1 and y2?

OpenStudy (archaists):

@OswaldMurphy It's because when you plot the points and then calculate the slope, it will give you 4. Then you can plug it into y=mx+b, to find b, which will be -6. Function notation just means that it will have f(x).

OpenStudy (oswaldmurphy):

Since I am finding x here do I replace both x's with 15? (another problem, don't solve just answer this please :D) Given that f(x) = 2x − 5, find the value of x that makes f(x) = 15

OpenStudy (archaists):

Is this another question?

OpenStudy (archaists):

Ohhhh, I see, x=20 then. f(x)=2x-5 f(x)=2*10-5 f(x)=20-5 f(x)=15